Week of 200907

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes



  • local:
  • remote: Miro (DB, Chair), Alberto (Monitoring), Borja (Monitoring), Darren (RAL), Elena (CNAF), Gavin (Compute), Xavier (KIT), Pavlo (ATLAS), Ville (NDGF), David (IN2P3), Maarten (ALICE), Mark (LHCb), Dave (FNAL), Onno (NL-T1)

Experiments round table:

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • Activity:
      • MC and WG productions; User jobs.
    • Issues:
      • CNAF: Issue discovered with Storm release from ~10 days ago that causes puppet to restart service. Puppet temporarily stopped until fix is released.
      • IN2P3: Issue with users unable to download data from storage due to auth problem (GGUS:148550)

Sites / Services round table:

  • ASGC: NC
  • BNL: NTR
  • EGI: NC
  • IN2P3: NTR
  • KIT: SRRs are now generated for all dCache SEs. Though accounting details for tape are updated daily, not every 30 minutes.
  • NDGF: dCache pool upgrades tomorrow starting at 09:00 CEST. It should not have much impact on anything.
  • NL-T1: Surfsara tape system is unstable. This may cause delays in restoring data. Vendor has been contacted for support.
  • NRC-KI: NC
  • OSG: NC
  • PIC: NTR
  • RAL: NTR

  • CERN computing services: NTR
  • CERN storage services: NC
  • CERN databases: NTR
  • Monitoring:
    • Draft reports for the August availability sent around
  • MW Officer: NC
  • Networks: NTR
  • Security: NTR


Edit | Attach | Watch | Print version | History: r17 < r16 < r15 < r14 < r13 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r17 - 2020-09-08 - NikolayVoytishin
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback